OpenStudy (anonymous):

Determine whether the function is periodic. If it is periodic, find the period. f(x) = 3 sin 2x + 4 cos 3x a)2pi b)6pi c)not periodic d)pi @phi @mathstudent55 @thomaster @Catlover5925 @Kainui

OpenStudy (catlover5925):

ummm sorry I have never done this

OpenStudy (kainui):

If it's periodic then it repeats after some time so we can plug it in like this: \[\LARGE f(x)=f(x+T)\] we already know that sine and cosine are periodic themselves so we can use this fact to help us find the period of f(x) \[\LARGE \sin(x)=\sin(x+2\pi)\]